# Ultralytics YOLO Evolution: An Overview of YOLO26, YOLO11, YOLOv8, and YOLOv5 Object Detectors for Computer Vision and Pattern Recognition

### 15 Oct 2025 v2

@@ 1 Introduction
Object detection has emerged as one of the most critical tasks in computer vision, enabling machines not only to identify but also to localize multiple objects within complex image or video streams zhao2019object ; zou2023object . Its importance spans a wide spectrum of domains, including autonomous driving, robotics, surveillance, medical imaging, agriculture, and smart manufacturing, where reliable real-time performance directly translates into safety, efficiency, and automation gains rana2024artificial ; khan2025objectdetection . Among the numerous algorithms proposed over the past decade, the You Only Look Once (YOLO) family has emerged as the most influential and widely adopted series of models for real-time object detection, striking a balance between high accuracy and unprecedented inference speed sapkota2025yolo . Since its initial release in 2016, YOLO has undergone multiple architectural revisions, each iteration addressing specific limitations of its predecessors while incorporating advances in neural network design, training strategies, loss functions, and deployment efficiency sapkota2025yolo . The cumulative evolution of YOLO is visualized in Figure 1, which traces the transition from early Ultralytics releases such as YOLOv5 to the latest YOLO26 (2025), the first version to natively unify five key tasks: object detection, instance segmentation, classification, pose/keypoints detection, and oriented bounding box detection. This timeline illustrates how YOLO has steadily expanded its capabilities beyond detection to become a versatile multi-task vision framework suitable for both research and edge deployment.
The release of YOLO26 in September 2025 represents the latest milestone in this evolutionary trajectory. Designed around principles of simplicity, efficiency, and innovation, YOLO26 introduces architectural simplifications that remove bottlenecks such as Non-Maximum Suppression (NMS) and Distribution Focal Loss (DFL), adopts a novel MuSGD optimizer for stable convergence, and incorporates training refinements such as Progressive Loss Balancing (ProgLoss) and Small-Target-Aware Label Assignment (STAL) sapkota2025yolo26 . Together, these innovations establish YOLO26 as a cutting-edge model optimized for low-power and embedded devices, significantly improving real-world deployability without compromising accuracy.
@@
To situate YOLO26 within the broader ecosystem, we first recall the paradigm shift inaugurated by Redmon et al. (2016), who reframed detection as a single regression problem eschewing region proposals in favor of direct, one-shot prediction of bounding boxes and class probabilities redmon2016you . In contrast to two-stage pipelines such as R-CNN and Faster R-CNN, which decouple proposal generation and classification he2017mask ; ren2016faster , the YOLO design achieved real-time throughput while preserving competitive accuracy, making YOLOv1 attractive for latency-critical domains including robotics and autonomous navigation ali2024yolo ; diwan2023object . Rapid iterations followed: YOLOv2 (2017) introduced batch normalization, anchor boxes, and multi-scale training for robustness across object scales redmon2017yolo9000 ; YOLOv3 (2018) deepened the backbone (Darknet-53) and exploited multi-scale feature maps, establishing a widely adopted baseline for both academia and industry redmon2018yolov3 ; apostolidis2025delving ; sapkota2025yolo ; edozie2025comprehensive . As accuracy demands grew across aerial imaging, agriculture, and medical analysis, architectures diversified. YOLOv4 (2020) incorporated CSPNet, improved activations, and advanced training strategies (e.g., mosaic augmentation, CIoU) bochkovskiy2020yolov4 . Ultralytics’ YOLOv5 (2020) popularized a PyTorch-native, modular toolchain that eased adaptation to segmentation, classification, and edge deployment. Subsequent community releases (YOLOv6, YOLOv7) integrated parameter-efficient modules and transformer-inspired blocks to push accuracy while maintaining real-time inference li2022yolov6 ; wang2023yolov7 . Ultralytics then re-architected the stack with YOLOv8 (decoupled head, anchor-free predictions) sohan2024review ; farooq2024improved , followed by YOLOv9 (GELAN, progressive distillation) wang2024yolov9 , YOLOv10 (latency-balanced assignments) wang2024yolov10 , and YOLO11 (efficiency with strong small-object performance) sapkota2025yolo . Parallel non-Ultralytics lines YOLOv12 and YOLOv13 pursued attention-centric designs (multi-head self-attention, improved fusion, stronger regularization) but continued to rely on NMS and DFL, creating latency and export friction on low-power devices tian2025yolov12 ; lei2025yolov13 . Addressing these bottlenecks, YOLO26 advances a deployment-first philosophy: it removes DFL, adopts end-to-end (NMS-free) inference, and introduces ProgLoss and STAL for stability and small-object fidelity, while training is accelerated via MuSGD. Figure 1 summarizes this trajectory and clarifies YOLO26 as the first Ultralytics release that natively unifies object detection, instance segmentation, classification, pose/keypoints detection, and oriented bounding box detection.
@@ 2 Architectural Evolution of Ultralytics YOLO Models
To provide a precise chronological and architectural context, we separate Ultralytics-maintained releases from community-driven variants. Table 1 catalogs the Ultralytics lineage YOLOv5 (2020), YOLOv8 (2023), YOLO11 (2024), and YOLO26 (2025) highlighting design choices, capabilities, and deployment characteristics within a single vendor family. In contrast, Table 2 surveys major community-driven releases independent of Ultralytics, including YOLOv1 (2015), YOLOv2 (2016), YOLOv3 (2018), YOLOv4 (2020), YOLOv6 (2022), YOLOv7 (2022), YOLOv9 (2024), and YOLOv10 (2024), alongside other contemporary variants. Together, these tables present a coherent, year-by-year view of architectural innovations, task expansions, and performance trends across the YOLO ecosystem, clarifying how the framework matured into the state represented by YOLO26 (Tables 12).
@@
YOLOv1 (2015) redmon2016you First unified single-stage detector; predicted bounding boxes + class probabilities in a single forward pass, pioneering real-time object detection. Object Detection, Classification Darknet
YOLOv2 (2016) redmon2017yolo9000 Introduced anchor boxes and dimension clustering; enabled multi-scale training; YOLO9000 demonstrated joint detection and classification across 9000 categories. Object Detection, Classification Darknet
YOLOv3 (2018) redmon2018yolov3 Deeper Darknet-53 backbone with residual blocks; multi-scale predictions enhanced small-object detection; added SPP for richer feature context. Object Detection, Multi-scale Detection Darknet
YOLOv4 (2020) bochkovskiy2020yolov4 CSPDarknet-53 backbone and Mish activation; improved training stability; optimized for high accuracy-speed tradeoff on GPUs. Object Detection, Object Tracking Darknet
YOLOv6 (2022) li2022yolov6 Industrial-focused release (by Meituan): EfficientRep backbone, improved self-attention, anchor-free head; focused on deployment efficiency for production. Object Detection, Instance Segmentation PyTorch
YOLOv7 (2022) wang2023yolov7 Introduced E-ELAN for extended feature learning; re-parameterized convolution modules; incorporated transformer blocks for context modeling; optimized for tracking and multi-task vision. Object Detection, Tracking, Segmentation PyTorch
YOLOv9 (2024) wang2024yolov9 Introduced Programmable Gradient Information (PGI) for enhanced learning and G-ELAN backbone; boosted training flexibility and feature richness. Object Detection, Instance Segmentation PyTorch
YOLOv10 (2024) wang2024yolov10 Achieved end-to-end NMS-free detection using consistent dual-assignment training strategy; removed reliance on post-processing. Object Detection PyTorch
YOLOv12 (2025) tian2025yolov12 Attention-centric design: efficient area attention module and Residual ELAN (R-ELAN) blocks; achieved transformer-level accuracy at YOLO speed. Object Detection PyTorch
YOLOv13 (2025) lei2025yolov13 Proposed HyperACE (hypergraph-based correlation enhancement) and FullPAD (pipeline-wide aggregation/distribution) for global high-order feature learning; depthwise separable convolutions improved efficiency. Object Detection PyTorch
@@ 2.1 YOLO26: DFL Removal, NMS-Free Inference, ProgLoss, STAL, MuSGD
Figure 2: YOLO26 simplified architecture diagram
@@ 2.3 YOLOv8: Decoupled Head, Anchor-Free Predictions
YOLOv8 delivered a more principled detector head and a modernized representation pipeline. The decoupled head separates classification and regression branches, mitigating gradient interference between localization and recognition objectives fang2024lightweight . This design reduces optimization entanglement improving convergence smoothness and yielding better calibration at a given operating point (e.g., at fixed mAP or recall). Complementing this head, YOLOv8 formalized an anchor-free assignment strategy: instead of regressing offsets around predefined anchor priors, the model predicts normalized box parameters and confidences directly for candidate points on the feature map fang2024lightweight ; liang2025vst ; dolatyabi2025deep . Anchor-free heads simplify hyperparameter tuning (no anchor clustering) and improve generalization across datasets with differing object aspect-ratio statistics liu2024using ; zhang2024dsp .
@@
The backbone/neck stack evolved toward lighter, high-throughput modules (e.g., C2f-style blocks) that maintain receptive-field richness while reducing memory bandwidth pressure wang2023bl ; yang2023lightweight . Multi-scale features are fused with attention to stride alignment and aliasing minimization, preserving crisp spatial information critical to small objects. Training refinements included improved label assignment heuristics, class-imbalance mitigation, and stronger regularization under aggressive augmentation, which together stabilized long-horizon training and reduced sensitivity to batch size. These choices enabled coherent scaling across nano to extra-large variants without pathological underfitting or collapse at the smallest scales.
@@ 2.4 YOLOv5: Modularity and PyTorch Adoption
Ultralytics’ YOLOv5 marked a decisive shift from Darknet to a fully PyTorch-native implementation, catalyzing adoption by unifying training utilities, data pipelines, augmentation strategies, and export tooling under a single, accessible codebase jocher2022ultralytics . Architecturally, YOLOv5 popularized a family of depth/width–scaled variants (n/s/m/l/x) built on CSP-inspired backbones and PANet-style necks to improve multi-scale aggregation with minimal latency overhead ren2022yolov5 ; dang2023accelerated . The transition to SiLU activations and careful normalization choices reduced optimization pathologies (e.g., vanishing gradients in deeper stacks) while preserving high-frequency detail important for small-object detection jocher2021ultralytics .
A key contribution of YOLOv5 is modularity: components (backbone blocks, neck layers, heads, losses) can be swapped, ablated, or extended with minimal refactoring johnson2025yolo . This modularity underpins rapid task adaptation e.g., adding a segmentation head or a classification head without destabilizing the base detector. On the training side, YOLOv5 systematized strong, production-friendly augmentations (mosaic, copy-paste, HSV jitter, random perspective) and reliable schedules (cosine, one-cycle), paired with EMA weight tracking for robust final checkpoints. The data loader and caching logic were engineered for throughput on commodity hardware, enabling high iteration rates that accelerate experimentation cycles.
From a deployment perspective, YOLOv5 leaned into broad format support (ONNX, TensorRT, CoreML), mixed-precision inference (FP16), and lightweight post-processing nnadozie2024simplifying ; fang2021deployment ; yan2021real . Although it retained NMS and conventional bounding-box regression (with optional DFL in later forks), the streamlined graph and conservative operator set made it particularly amenable to embedded and cloud accelerators. Collectively, these design choices positioned YOLOv5 as a practical baseline: not the most novel architecturally, but extremely effective in closing the gap between research-grade code and production-grade inference at scale.
@@ 4.1 Export formats and hardware compatibility
Modern YOLO implementations emphasise portability munteanu2022sea ; ali2024yolo ; ray2025llmyoloedge . The Ultralytics ecosystem provides export scripts that convert PyTorch checkpoints into multiple deployment formats, enabling efficient inference on diverse hardware jocher2022ultralytics . YOLOv5 introduced a comprehensive export pipeline: models can be saved as native PyTorch weights (.pt), TorchScript (.torchscript) for embedding in C++ applications, ONNX (.onnx) for cross‑framework interoperability, OpenVINO for Intel CPUs and VPUs, TensorRT (.engine) for NVIDIA GPUs, CoreML (.mlmodel) for iOS devices, TensorFlow SavedModel, GraphDef, and TFLite (.tflite) for Android and microcontroller deployment, as well as specialised formats for Edge TPUs, TensorFlow.js, and PaddlePaddle jocher2020ultralytics ; kim2022object . Export to ONNX or OpenVINO typically yields up to three‑fold acceleration on CPUs, while conversion to TensorRT can deliver five‑fold GPU speedups. Table 5 summarises the major export options and the hardware targets they enable for each Ultralytics release.
@@
YOLOv8 refined the export interface, offering command‑line flags to select formats and optimise for mixed precision (e.g. FP16). YOLO11 retained this flexibility while adding orientated bounding box support in ONNX and TensorRT exports sapkota2025yolo . YOLO26’s simplified head and DFL removal further streamline conversion: without distributional regression or NMS, exported graphs contain fewer custom operations and thus map more cleanly to ONNX, TensorRT, CoreML, and TFLite backends. This improved portability ensures models trained in PyTorch can be embedded into C++ pipelines, mobile applications, or web environments with minimal engineering overhead.

### 16 Mar 2026 v3

@@ 1 Introduction
Object detection has emerged as one of the most critical tasks in computer vision, enabling machines not only to identify but also to localize multiple objects within complex image or video streams Zhao et al. (2019); Zou et al. (2023). Its importance spans a wide spectrum of domains, including autonomous driving, robotics, surveillance, medical imaging, agriculture, and smart manufacturing, where reliable real-time performance directly translates into safety, efficiency, and automation gains Rana and others (2024); Khan et al. (2025). Among the numerous algorithms proposed over the past decade, the You Only Look Once (YOLO) family has emerged as the most influential and widely adopted series of models for real-time object detection, striking a balance between high accuracy and unprecedented inference speed Sapkota et al. (2025b). Since its initial release in 2016, YOLO has undergone multiple architectural revisions, each iteration addressing specific limitations of its predecessors while incorporating advances in neural network design, training strategies, loss functions, and deployment efficiency Sapkota et al. (2025b). The cumulative evolution of YOLO is visualized in Figure 1, which traces the transition from early Ultralytics releases such as YOLOv5 to the latest YOLO26 (2025), the first version to natively unify five key tasks: object detection, instance segmentation, classification, pose/keypoints detection, and oriented bounding box detection. This timeline illustrates how YOLO has steadily expanded its capabilities beyond detection to become a versatile multi-task vision framework suitable for both research and edge deployment.
The release of YOLO26 in September 2025 represents the latest milestone in this evolutionary trajectory. Designed around principles of simplicity, efficiency, and innovation, YOLO26 introduces architectural simplifications that remove bottlenecks such as Non-Maximum Suppression (NMS) and Distribution Focal Loss (DFL), adopts a novel MuSGD optimizer for stable convergence, and incorporates training refinements such as Progressive Loss Balancing (ProgLoss) and Small-Target-Aware Label Assignment (STAL) Sapkota et al. (2025a). Together, these innovations establish YOLO26 as a cutting-edge model optimized for low-power and embedded devices, significantly improving real-world deployability without compromising accuracy.
@@
To situate YOLO26 within the broader ecosystem, we first recall the paradigm shift inaugurated by Redmon et al. (2016), who reframed detection as a single regression problem eschewing region proposals in favor of direct, one-shot prediction of bounding boxes and class probabilities Redmon et al. (2016). In contrast to two-stage pipelines such as R-CNN and Faster R-CNN, which decouple proposal generation and classification He et al. (2017); Ren et al. (2016), the YOLO design achieved real-time throughput while preserving competitive accuracy, making YOLOv1 attractive for latency-critical domains including robotics and autonomous navigation Ali and Zhang (2024); Diwan et al. (2023). Rapid iterations followed: YOLOv2 (2017) introduced batch normalization, anchor boxes, and multi-scale training for robustness across object scales Redmon and Farhadi (2017); YOLOv3 (2018) deepened the backbone (Darknet-53) and exploited multi-scale feature maps, establishing a widely adopted baseline for both academia and industry Redmon and Farhadi (2018); Apostolidis and Papakostas (2025); Sapkota et al. (2025b); Edozie et al. (2025). As accuracy demands grew across aerial imaging, agriculture, and medical analysis, architectures diversified. YOLOv4 (2020) incorporated CSPNet, improved activations, and advanced training strategies (e.g., mosaic augmentation, CIoU) Bochkovskiy et al. (2020). Ultralytics’ YOLOv5 (2020) popularized a PyTorch-native, modular toolchain that eased adaptation to segmentation, classification, and edge deployment. Subsequent community releases (YOLOv6, YOLOv7) integrated parameter-efficient modules and transformer-inspired blocks to push accuracy while maintaining real-time inference Li et al. (2022a); Wang et al. (2023a). Ultralytics then re-architected the stack with YOLOv8 (decoupled head, anchor-free predictions) Sohan et al. (2024); Farooq et al. (2024), followed by YOLOv9 (GELAN, progressive distillation) Wang et al. (2024b), YOLOv10 (latency-balanced assignments) Wang et al. (2024a), and YOLO11 (efficiency with strong small-object performance) Sapkota et al. (2025b). Parallel non-Ultralytics lines YOLOv12 and YOLOv13 pursued attention-centric designs (multi-head self-attention, improved fusion, stronger regularization) but continued to rely on NMS and DFL, creating latency and export friction on low-power devices Tian et al. (2025); Lei et al. (2025). Addressing these bottlenecks, YOLO26 advances a deployment-first philosophy: it removes DFL, adopts end-to-end (NMS-free) inference, and introduces ProgLoss and STAL for stability and small-object fidelity, while training is accelerated via MuSGD. Figure 1 summarizes this trajectory and clarifies YOLO26 as the first Ultralytics release that natively unifies object detection, instance segmentation, classification, pose/keypoints detection, and oriented bounding box detection.
@@ 2 Architectural Evolution of Ultralytics YOLO Models
To provide a precise chronological and architectural context, we separate Ultralytics-maintained releases from community-driven variants. Table 1 catalogs the Ultralytics lineage YOLOv5 (2020), YOLOv8 (2023), YOLO11 (2024), and YOLO26 (2025) highlighting design choices, capabilities, and deployment characteristics within a single vendor family. In contrast, Table 2 surveys major community-driven releases independent of Ultralytics, including YOLOv1 (2015), YOLOv2 (2016), YOLOv3 (2018), YOLOv4 (2020), YOLOv6 (2022), YOLOv7 (2022), YOLOv9 (2024), and YOLOv10 (2024), alongside other contemporary variants. Together, these tables present a coherent, year-by-year view of architectural innovations, task expansions, and performance trends across the YOLO ecosystem, clarifying how the framework matured into the state represented by YOLO26 (Tables 1-2).
@@
YOLOv1 (2015) Redmon et al. (2016) First unified single-stage detector; predicted bounding boxes + class probabilities in a single forward pass, pioneering real-time object detection. Object Detection, Classification Darknet
YOLOv2 (2016) Redmon and Farhadi (2017) Introduced anchor boxes and dimension clustering; enabled multi-scale training; YOLO9000 demonstrated joint detection and classification across 9000 categories. Object Detection, Classification Darknet
YOLOv3 (2018) Redmon and Farhadi (2018) Deeper Darknet-53 backbone with residual blocks; multi-scale predictions enhanced small-object detection; added SPP for richer feature context. Object Detection, Multi-scale Detection Darknet
YOLOv4 (2020) Bochkovskiy et al. (2020) CSPDarknet-53 backbone and Mish activation; improved training stability; optimized for high accuracy-speed tradeoff on GPUs. Object Detection, Object Tracking Darknet
YOLOv6 (2022) Li et al. (2022a) Industrial-focused release (by Meituan): EfficientRep backbone, improved self-attention, anchor-free head; focused on deployment efficiency for production. Object Detection, Instance Segmentation PyTorch
YOLOv7 (2022) Wang et al. (2023a) Introduced E-ELAN for extended feature learning; re-parameterized convolution modules; incorporated transformer blocks for context modeling; optimized for tracking and multi-task vision. Object Detection, Tracking, Segmentation PyTorch
YOLOv9 (2024) Wang et al. (2024b) Introduced Programmable Gradient Information (PGI) for enhanced learning and G-ELAN backbone; boosted training flexibility and feature richness. Object Detection, Instance Segmentation PyTorch
YOLOv10 (2024) Wang et al. (2024a) Achieved end-to-end NMS-free detection using consistent dual-assignment training strategy; removed reliance on post-processing. Object Detection PyTorch
YOLOv12 (2025) Tian et al. (2025) Attention-centric design: efficient area attention module and Residual ELAN (R-ELAN) blocks; achieved transformer-level accuracy at YOLO speed. Object Detection PyTorch
YOLOv13 (2025) Lei et al. (2025) Proposed HyperACE (hypergraph-based correlation enhancement) and FullPAD (pipeline-wide aggregation/distribution) for global high-order feature learning; depthwise separable convolutions improved efficiency. Object Detection PyTorch
@@ 2.1 YOLO26: DFL Removal, NMS-Free Inference, ProgLoss, STAL, MuSGD
Figure 2: YOLO26 core architecture diagram
@@ 2.3 YOLOv8: Decoupled Head, Anchor-Free Predictions
YOLOv8 delivered a more principled detector head and a modernized representation pipeline. The decoupled head separates classification and regression branches, mitigating gradient interference between localization and recognition objectives Fang and Yang (2024). This design reduces optimization entanglement improving convergence smoothness and yielding better calibration at a given operating point (e.g., at fixed mAP or recall). Complementing this head, YOLOv8 formalized an anchor-free assignment strategy: instead of regressing offsets around predefined anchor priors, the model predicts normalized box parameters and confidences directly for candidate points on the feature map Fang and Yang (2024); Liang and Chen (2025); Dolatyabi et al. (2025). Anchor-free heads simplify hyperparameter tuning (no anchor clustering) and improve generalization across datasets with differing object aspect-ratio statistics Liu et al. (2024); Zhang et al. (2024).
@@
The backbone/neck stack evolved toward lighter, high-throughput modules (e.g., C2f-style blocks) that maintain receptive-field richness while reducing memory bandwidth pressure Wang et al. (2023b); Yang et al. (2023). Multi-scale features are fused with attention to stride alignment and aliasing minimization, preserving crisp spatial information critical to small objects. Training refinements included improved label assignment heuristics, class-imbalance mitigation, and stronger regularization under aggressive augmentation, which together stabilized long-horizon training and reduced sensitivity to batch size. These choices enabled coherent scaling across nano to extra-large variants without pathological underfitting or collapse at the smallest scales.
@@ 2.4 YOLOv5: Modularity and PyTorch Adoption
Ultralytics’ YOLOv5 marked a decisive shift from Darknet to a fully PyTorch-native implementation, catalyzing adoption by unifying training utilities, data pipelines, augmentation strategies, and export tooling under a single, accessible codebase Jocher et al. (2022). Architecturally, YOLOv5 popularized a family of depth/width–scaled variants (n/s/m/l/x) built on CSP-inspired backbones and PANet-style necks to improve multi-scale aggregation with minimal latency overhead Ren et al. (2022); Dang et al. (2023). The transition to SiLU activations and careful normalization choices reduced optimization pathologies (e.g., vanishing gradients in deeper stacks) while preserving high-frequency detail important for small-object detection Jocher et al. (2021).
A key contribution of YOLOv5 is modularity: components (backbone blocks, neck layers, heads, losses) can be swapped, ablated, or extended with minimal refactoring Johnson (2025). This modularity underpins rapid task adaptation e.g., adding a segmentation head or a classification head without destabilizing the base detector. On the training side, YOLOv5 systematized strong, production-friendly augmentations (mosaic, copy-paste, HSV jitter, random perspective) and reliable schedules (cosine, one-cycle), paired with EMA weight tracking for robust final checkpoints. The data loader and caching logic were engineered for throughput on commodity hardware, enabling high iteration rates that accelerate experimentation cycles.
From a deployment perspective, YOLOv5 leaned into broad format support (ONNX, TensorRT, CoreML), mixed-precision inference (FP16), and lightweight post-processing Nnadozie et al. (2024); Fang et al. (2021); Yan et al. (2021). Although it retained NMS and conventional bounding-box regression (with optional DFL in later forks), the streamlined graph and conservative operator set made it particularly amenable to embedded and cloud accelerators. Collectively, these design choices positioned YOLOv5 as a practical baseline: not the most novel architecturally, but extremely effective in closing the gap between research-grade code and production-grade inference at scale.
@@ 4.1 Export formats and hardware compatibility
Modern YOLO implementations emphasise portability Munteanu et al. (2022); Ali and Zhang (2024); Ray et al. (2025). The Ultralytics ecosystem provides export scripts that convert PyTorch checkpoints into multiple deployment formats, enabling efficient inference on diverse hardware Jocher et al. (2022). YOLOv5 introduced a comprehensive export pipeline: models can be saved as native PyTorch weights (.pt), TorchScript (.torchscript) for embedding in C++ applications, ONNX (.onnx) for cross‑framework interoperability, OpenVINO for Intel CPUs and VPUs, TensorRT (.engine) for NVIDIA GPUs, CoreML (.mlmodel) for iOS devices, TensorFlow SavedModel, GraphDef, and TFLite (.tflite) for Android and microcontroller deployment, as well as specialised formats for Edge TPUs, TensorFlow.js, and PaddlePaddle Jocher et al. (2020); Kim et al. (2022). Export to ONNX or OpenVINO typically yields up to three‑fold acceleration on CPUs, while conversion to TensorRT can deliver five‑fold GPU speedups. Table 5 summarises the major export options and the hardware targets they enable for each Ultralytics release.
@@
YOLOv8 refined the export interface, offering command‑line flags to select formats and optimise for mixed precision (e.g. FP16). YOLO11 retained this flexibility while adding orientated bounding box support in ONNX and TensorRT exports Sapkota et al. (2025b). YOLO26’s simplified head and DFL removal further streamline conversion: without distributional regression or NMS, exported graphs contain fewer custom operations and thus map more cleanly to ONNX, TensorRT, CoreML, and TFLite backends. This improved portability ensures models trained in PyTorch can be embedded into C++ pipelines, mobile applications, or web environments with minimal engineering overhead.